# Physics-Driven Local-Whole Elastic Deformation Modeling for Point Cloud Representation Learning

### v2

@@ III-B Physical Information Awareness
where $\nabla\cdot{\sigma_{il}}$} represents the divergence of the stress tensor, which describes the force equilibrium within the mesh. $f_{il}$ denotes the ground-truth external force at the $l$-th tetrahedron of the $i$-th sample. This loss term evaluates whether each tetrahedral element satisfies physical equilibrium, thereby serving as a physics-based constraint to guide the network toward generating reasonable predictions that comply with mechanical laws.
III-C Overall Object
@@ IV-A Pre-training
Baselines. We compare our method with the current mainstream self-supervised learning methods: Jigsaw3D [31] and Rotation [36] are based on pretext tasks. CPG[51], PointMoment[35], PointSmile [17], PoCCA [33], CLR-GAM [34], ToThePoint [37], and STRL [15] are based on contrast learning methods. TCMSS [39], HyperIPC[40], MCIB[41], CrossNet [16], Cosine Mixup [42], PointMCD [43], and CrossPoint [44] are based on cross-modal methods. GDANet[52], GSPCon [11], Tsai et al. [29], and OcCo [26] are based on the generative models. Additionally, for clarity in the tables below, we use the abbreviations Sup, PT, CL, CM, and GM to represent supervised learning, pretext tasks, contrastive learning, cross-modal methods, and generative models, respectively.
@@
Implementation Details. To facilitate a fair comparison with existing self-supervised learning methods, we follow the experimental setup proposed in [15] and use the backbone networks of PointNet [5] and DGCNN [7] as the encoder ${\Phi}_{\theta}$}. We use Adam [58] optimizer with weight decay $1\times{{10}^{-4}}$ and initial learning rate $1\times{{10}^{-3}}$. The learning rate scheduler is cosine annealing [59], and the model is trained end-to-end across 200 epochs. After pre-training, all downstream tasks are performed on the pre-trained encoder ${\Phi}_{\theta}$}. Our models are trained on two Tesla V100-PCI-E-32G GPUs. For the classification task, we use a pre-trained backbone model to extract point cloud features on the ModelNet40 dataset and ScanObjectNN dataset, and these features are used to train a simple linear Support Vector Machine (SVM) classifier [60]. For segmentation tasks, We evaluate part segmentation using the mean intersectionover-union (mIoU) metric, which is computed by averaging the IoU of each part in the object and then averaging the values obtained from each object category. In dataset S3DIS, we performed 6-fold cross-validation on its 6 areas to obtain the final segmentation results.
@@
Figure 4: We use t-SNE to visualise features extracted from PointNet and DGCNN on the ModelNet10 test set. Different color represents different categories.
@@ IV-B Object Classification
We evaluate the trained pre-trained weights on the ModelNet10 dataset by extracting features for each sample and performing t-SNE [61] clustering visualization, as shown in Figure 4. The left and right sides of the figure visualise the PointNet and DGCNN model features, respectively. It can be observed that our pre-trained model effectively distinguishes most samples according to their categories, resulting in clear classification boundaries.
@@ IV-C Object Segmentation
Figure 7: Visualization of elastic deformation under external forces across object categories. The first row shows the original shapes of the objects, while the second row presents the deformation results predicted by the network. The red arrows indicate the direction of the applied force, while the green dashed circles highlight regions exhibiting significant deformation.
@@
IV-D Ablation Study
Module Analysis. we conducted a qualitative analysis of the performance impact of each part on the ModelNet40 and ScanObjectNN datasets. Similar to the classification experiments, we performed pre-training for each part separately, followed by training an SVM classifier using the extracted point cloud features, as summarized in Table IV. The results demonstrate that incorporating physical information as a constraint into the model learning process enhances the detection performance, enabling the model to learn more discriminative point cloud representations.
@@ IV-D Ablation Study
[…]
TABLE VI: The time and memory required for tetrahedral mesh generation and deformation calculations.
Type time/s memory/MB
[…]
Deformation visualization. Figure 7 presents the predicted geometric deformations of various object categories under downward external force. The top row shows the original shapes in the undeformed state, while the bottom row illustrates the network-predicted elastic deformations. It can be observed that significant local deformations occur in structurally critical regions, such as the handle of a cup, the brim edge of a hat, the grip of a pistol, the joint of a lamp arm, and the load-bearing surface of a chair. These areas are typically stress-sensitive zones where force transmission leads to stress concentration. In the figure, these regions with pronounced deformation effects are highlighted using green dashed circles. Importantly, these local deformations do not occur in isolation but emerge progressively under the influence of global force propagation through the structure, reflecting the network’s ability to model the physical coupling between local parts and the global shape. The experimental results validate the effectiveness and soundness of the proposed physics-driven modeling framework in simulating the deformation process from localized responses to holistic structural reconstruction, demonstrating its capacity to capture the internal force transmission patterns within complex objects.
@@ V Conclusion
This paper proposes a physics-driven self-supervised learning method for point cloud representation, which effectively models structural responses under force propagation by introducing a local-whole deformation mechanism derived from elasticity theory. The framework consists of two training modules: one models the whole geometric structure using an implicit field, enabling expressive representation for arbitrary shapes; the other captures elastic responses induced by external forces based on tetrahedral mesh construction and constrains deformation behavior through a physics loss. Both branches share encoder weights, establishing a closed-loop mapping between local force-induced deformations and global shape changes, thereby deeply integrating local details with global structural features. Experimental results demonstrate that the proposed method achieves superior performance across multiple downstream tasks, including classification and segmentation, consistently outperforming existing state-of-the-art approaches. These results validate the effectiveness of incorporating physics-based modeling for enhancing the robustness and generalization capability of point cloud representations.
Nevertheless, the current method still has limitations. For instance, in complex non-rigid deformation scenarios, there remains room for improving the physical accuracy of tetrahedral mesh construction and deformation simulation. Additionally, how to further integrate force propagation modeling with object properties in real-world environments remains an important direction for future research. Subsequent work will focus on more efficient physical simulation techniques to advance point cloud understanding from purely geometric representations toward deeper integration with structural and physical attributes, enabling more interpretable and generalizable modeling of real-world objects.

### v3

@@ III-C Physical Information Awareness
where $\nabla\cdot{\sigma_{il}}$} represents the stress tensor divergence, which describes the force equilibrium within the mesh. $f_{il}$ denotes the ground-truth external force at the $l$-th tetrahedron of the $i$-th sample. This loss term evaluates whether each tetrahedral element satisfies physical equilibrium, thereby serving as a physics-based constraint to guide the network toward generating reasonable predictions that comply with mechanical laws.
III-D Overall Object
@@ IV-A Pre-training
Baselines. We compare our method with the current mainstream self-supervised learning methods: Jigsaw3D [33] and Rotation [37] are based on pretext tasks. CPG[56], CCPoint [39], PointSmile [17], PoCCA [35], CLR-GAM [36], ToThePoint [38], and STRL [15] are based on contrast learning methods. CrossCon-Jig [20], TCMSS [41], HyperIPC[42], MCIB[43], CrossNet [16], Cosine Mixup [44], PointMCD [45], and CrossPoint [46] are based on cross-modal methods. MAE3D [32], GDANet[57], GSPCon [11], Tsai et al. [30], and OcCo [27] are based on the generative models. Additionally, for clarity in the tables below, we use the abbreviations Sup, PT, CL, CM, and GM to represent supervised learning, pretext tasks, contrastive learning, cross-modal methods, and generative models, respectively.
@@
Implementation Details. To facilitate a fair comparison with existing self-supervised learning methods, we follow the experimental setup proposed in [15] and use the backbone networks of PointNet [5] or DGCNN [7] as the encoder ${\Phi}_{\theta}$}. We use Adam [63] optimizer with weight decay $1\times{{10}^{-4}}$ and initial learning rate $1\times{{10}^{-3}}$. The learning rate scheduler is cosine annealing [64], and the model is trained end-to-end across 200 epochs. After pre-training, all downstream tasks are performed on the pre-trained encoder ${\Phi}_{\theta}$}. Our models are trained on two Tesla V100-PCI-E-32G GPUs. For the classification task, we use a pre-trained backbone model to extract point cloud features on the ModelNet40 dataset and ScanObjectNN dataset, and these features are used to train a simple linear Support Vector Machine (SVM) classifier [65]. For segmentation tasks, We evaluate part segmentation using the mean intersectionover-union (mIoU) metric, which is computed by averaging the IoU of each part in the object and then averaging the values obtained from each object category. In dataset S3DIS, we performed 6-fold cross-validation on its 6 areas to obtain the final segmentation results. Additional experiments are provided in the supplementary material.
@@ IV-B Object Classification
Figure 4: Feature-space visualization (t-SNE) of learned point-cloud embeddings on ShapeNet with representative baselines. (a) Ours with PointNet backbone, (b) Ours with DGCNN backbone, (c) STRL (DGCNN), and (d) CrossPoint (DGCNN). Colors indicate object categories.
We evaluate the trained pre-trained weights on the ModelNet10 dataset by extracting features for each sample and performing t-SNE [66] clustering visualization, as shown in Figure 4. (a) and (b) show the feature distributions obtained by our method with PointNet and DGCNN backbones, respectively, while (c) and (d) correspond to two representative baselines: the contrastive learning method STRL (with DGCNN backbone) and the cross-modal method CrossPoint (with DGCNN backbone). As can be seen, our method forms more compact intra-class clusters and clearer inter-class separation under both backbones. Most categories exhibit well-defined clustering boundaries and fewer outliers, indicating more stable representational capabilities (the consistent trend from (a) to (b) further suggests that our method is relatively robust to the choice of backbone). In contrast, the feature distributions of STRL and CrossPoint tend to show more intra-class dispersion and elongated cluster shapes, suggesting that the representations they learn are less robust in terms of structural consistency. Our approach is capable of learning more structured point cloud representations.
@@
[…]
@@ IV-C Object Segmentation
IV-D Ablation Study
Module Analysis. we conducted a qualitative analysis of the performance impact of each part on the ModelNet40 and ScanObjectNN datasets. Following the same protocol as the classification experiments, we pre-train the backbone under three module settings: Physical Information Awareness (PIA), Implicit Feature Learning (IFL), and their combination (PIA+IFL). We then train a linear SVM on the extracted point-cloud features, as summarized in Table IV. The results demonstrate that incorporating physical information as a constraint into the model learning process enhances the detection performance, enabling the model to learn more discriminative point cloud representations.
Figure 7: Visualization of deformation and saliency feature. From top to bottom: (1) original shapes; (2) predicted elastic deformations under the same downward loading; (3) visualization of the load response (von Mises equivalent stress $\sigma_{VM}$} computed from the displacement field is rendered to characterize the spatial distribution of the load response, with colors varying from cool to warm indicating increasing response magnitude); (4)(5) Point-wise saliency heatmaps generated by the models without and with physical constraints, respectively, used to compare the differences in attended regions under the two settings (colors from cool to warm indicate increasing saliency). The red arrows indicate the direction of the applied force, while the green dashed circles highlight regions exhibiting significant deformation.
@@ IV-D Ablation Study
Deformation visualization and saliency analysis. To further provide an intuitive analysis of the model’s response patterns to deformation, we jointly visualize the predicted deformations and feature saliency. As shown in Figure 7, in the first three rows, the most pronounced local deformations are concentrated in structurally critical regions (e.g., mug handles, hat brims, pistol barrels, lamp arm joints, and load-bearing surfaces of chairs). These regions are typically stress-sensitive and therefore more prone to stress concentration (highlighted by green dashed circles). Importantly, these local deformations do not occur in isolation but arise as a result of under the influence of global force propagation through the structure, reflecting the network’s ability to model the physical coupling between local parts and the global shape. Comparing with the fourth row of Figure 7 (the variant without physical constraints), we observe that its saliency responses tend to be more scattered and are more easily affected by local geometric details or noise. In contrast, the saliency maps in the fifth row (with the physics-based mechanism enabled) are more strongly focused on the primary load-bearing regions and their neighboring load-transfer paths, and exhibit a much clearer spatial correspondence with the stress distribution. These results indicate that the physics-driven mechanism guides the network to shift its attribution from dispersed local geometric perturbations toward key load-bearing areas, thereby more faithfully capturing the propagation from local force responses to global deformations and improving the consistency and interpretability of the local-to-whole modeling.
Due to space limitations, additional ablation and analysis results are deferred to the supplementary material. Specifically, we include (i) qualitative module analysis; (ii) a hyperparameter study of the loss weights $a$ and $b$; (iii) implementation details and the physical configuration of the FEM-based module (parameterization and pseudocode); (iv) tetrahedral mesh generation and quality analysis, covering volume-based filtering as well as geometric fidelity and element quality for the Delaunay mesh; and (v) a training effectiveness analysis, reporting preprocessing overhead and training-time performance comparisons.
@@ V Conclusion
This paper proposes a physics-driven self-supervised learning method for point cloud representation, which effectively models structural responses under force propagation by introducing a local-whole deformation mechanism derived from elasticity theory. The framework consists of two training modules: one models the whole geometric structure using an implicit field, enabling expressive representation for arbitrary shapes; the other captures elastic responses induced by external forces based on tetrahedral mesh construction and constrains deformation behavior through a physics loss. Both branches share encoder weights, establishing a closed-loop mapping between local force-induced deformations and global shape changes, thereby deeply integrating local details with global structural features. Experimental results demonstrate that the proposed method achieves superior performance across multiple downstream tasks, including classification and segmentation. These results validate the effectiveness of incorporating physics-based modeling for enhancing the robustness and generalization capability of point cloud representations.